OpenStudy (anonymous):

The noodles & Company location in Parker, Colorado,sells its Pad Thai for $4.25 for a small,$5.25 for a regular with an additional $2.00 for protein for either size. If Noodles & Company is making a 300% markup based on the cost on their food,what is the cost for a small size Pad Thai with protein? Round to the nearest cent.

OpenStudy (anonymous):

$4.25 + $2.00 = $6.25 $6.25/3.0 = 2.08 6.25-2.08 = $4.17 I think that's right- give it a shot

OpenStudy (anonymous):

Okay jfcarlucci, but don't markup mean higher than it is already is

OpenStudy (anonymous):

I was thinking ($4.25+$2.00)=$6.25 $6.25*3.0=$18.75

OpenStudy (anonymous):

If not $18.75, it could possible be $19.00... $4.25*3.0+$4.25=$17.00+$2.00=$19.00 Any suggestions

OpenStudy (anonymous):

It doesn't. It would actually be $2.08. If your costs are $6.25 and that is a markup of 300% the cost of that item without the markup is $2.08.

OpenStudy (anonymous):

So jfcarlucci with this problem, the markup based on the cost was 300%. What is the corresponding markup based on the selling price?